Find the Wronskian of the following functions and determine whether it is linearly dependent or linearly independent on (-∞,∞).

  1. {ln x, ln x2}                   ans, W=0, linearly dependent
  2. {2+x, 1-x, 3+x2}           ans, W=-6, linearly independent 
1
Expert's answer
2021-12-03T11:06:13-0500

1.

"\{\ln x, \ln x^2\}"

"\{\ln x, 2\ln x\}, x>0"

"(\ln x)'=1/x, (\ln(x^2))'=2/x"

"W(\ln x, \ln x^2)=\begin{vmatrix}
 \ln x & \ln x^2 \\
 1/x & 2/x
\end{vmatrix}"

"=(2/x)\ln x-(1/x)\ln x^2"

"=(2/x)\ln x-(2/x)\ln x=0"

"\ln x, \ln x^2" are linearly dependent on "(0, \infin)."


2.


"\{2+x, 1-x, 3+x^2\}"

"(2+x)'=1,(1-x)'=-1, (3+x^2)'=2x"

"(2+x)''=0,(1-x)''=0, (3+x^2)''=2"

"W(2+x, 1-x,3+x^2)=\begin{vmatrix}
 2+x & 1-x & 3+x^2 \\
 1 & -1 & 2x \\
 0 & 0 & 2
\end{vmatrix}"


"=2\begin{vmatrix}
 2+x & 1-x \\
 1 & -1
\end{vmatrix}=2(-2-x-1+x)"

"=-6\not=0"

"2+x, 1-x, 3+x^2" are linearly independent on "(-\infin, \infin)."